# Learning Robust Social Strategies with Large Language Models

### v1

@@ 5.1 Naive MARL leads to greedy behavior with LLMs
Figure 2: Training curves of multi-agent RLOO on several open-source LLMs across IPD, Split No-Comm, and Trust and Split. In all environments, average rewards converge to the greedy payoff levels, showing that naive MARL drives LLMs toward defecting strategies in social dilemmas.
In order to robustly demonstrate how MARL interacts with LLMs in social dilemmas, we train LLMs from several model families across all the environments. We use multi-agent RLOO with self-play as the learning algorithm and train only the LoRA parameters. Figure 2 shows that naive MARL consistently converges to greedy behavior across all environments and model families. In simpler environments like IPD, all models begin with higher than greedy average rewards but drift toward greedy play with training. In more complex environments such as Split No-Comm and Trust and Split, Qwen models briefly achieve higher average rewards than greedy play before collapsing back to greedy behavior, while Llama and Gemma models start with low performance and converge directly to greedy strategies. Qualitatively, in Split No-Comm, we find that agents learned to bid the highest for every item even when they value it less. In Trust and Split, agents communicate their private hands honestly but in the proposals both always take all of coins for themselves. These results show that naive MARL robustly leads to greedy behavior in social dilemma settings. Since LLMs agents are likely to operate in scenarios that involve social dilemmas, this highlights the need for training methods that enable robust cooperation without being exploitable.
Next, we scale our experiments by training a Qwen-2.5-7B-Instruct agent against a frozen GPT-5-nano opponent using naive MARL. In this setup, Alice is the learning agent trained with LoRA, while Bob is the fixed GPT-5-nano. Figure 7 (left) shows that the RL-trained agent steadily exploits the fixed GPT-5-nano in Trust and Split: the average reward of the RL agent increases throughout training, while GPT-5-nano’s reward declines. Early in the training, the RL agent plays poorly and receives lower rewards, as it must learn how to act in the environment. After about 150 training steps, it begins exploiting GPT-5-nano, and the reward gap consistently widens. Finally, from the conversations shown in Figure 3, we observe that the RL-trained agent sometimes misrepresents the dominance relation between hands (for example, claiming that scissors is the upper hand over rock) and pairs this with a proposal giving itself a larger share. GPT-5-nano accepts these proposals, indicating that the fixed model is susceptible to strategically misleading communication. These results underscore that RL-trained agents become increasingly greedy in social dilemmas and that even state-of-the-art closed-source models remain vulnerable to such exploitation.
Figure 3: Example interaction in Trust and Split where an agent trained with multi-agent RLOO misrepresents the rock-paper-scissors hierarchy to claim the high value role. GPT-5 nano accepts the deceptive claim and proposal, illustrating that a fixed advanced model can be exploited by an RL-trained agent.
@@
Figure 4: Average rewards when evaluating an Advantage Alignment (AdAlign) agent, an always-cooperate (AC) agent, and an always-defect (AD) agent. In IPD (left) and Split No-Comm (right), Advantage Alignment achieves near cooperative payoffs with itself and AC while remaining robust against AD. Results are averaged over eight AdAlign agents trained with different random seeds
Figure 5: Average reward in Trust and Split when pitting an Advantage agent against agents trained with multi-agent RLOO (Defectors), and multi-agent RLOO with sum of rewards (Cooperators). Advantage Alignment cooperates with cooperative partners and itself, yet avoids being exploited by greedy agents. Results are averaged over eight AdAlign agents trained with different random seeds.
@@ 5.2 Advantage Alignment learns robust social strategies
To address the shortcomings of naive MARL, we apply Advantage Alignment to learn robust policies in our environments. We run Advantage Alignment with eight different seeds across all environment and report average results in Figure 4. For simpler environments such as IPD and Split No-Comm, the baseline agents, always-cooperate and always-defect agents can be hardcoded. In IPD, Always-Cooperate (AC) agent always plays action A, equivalent to Cooperate and the Always-Defect (AD) agent always plays action B equivalent to Defect as defined in section 2.3. In Split No-Comm, the AC agent proposes 10 when its own value is $10$ and the other player’s value is 1, proposes 0 in the reverse case, and proposes 5 when both values are equal. The AD agent always proposes 10, regardless of the values. In the Iterated Prisoner’s Dilemma, Advantage Alignment agents cooperate with themselves and with fully cooperative agents, while remaining robust against defection. The slight drop in performance against defectors comes from losing the first round, since the agent initially cooperates and receives a lower payoff on that round. In the Split No-Comm game, Advantage Alignment agents obtain about 86% of the full cooperation efficiency while still maintaining robustness. When paired with defectors, their performance decreases only slightly, indicating they are not easily exploitable. Qualitatively, Advantage Alignment learns a tit-for-tat strategy in IPD: it defects when the other agent defected in the previous round, and cooperates when the other agent cooperated. In Split No-Comm, it learns a strategy closer to grim-trigger, where a single defection can lead to persistent defection afterward.
In Trust and Split, we cannot hardcode cooperative and defective policies because the environment requires communication. Instead, we train baseline agents using multi-agent RLOO, and its sum-of-rewards variant. As shown in Figure 5, multi-agent RLOO produces defectors that achieve low average reward when paired with themselves, while the sum-of-rewards variant produces cooperators that achieve the maximum possible reward with themselves. However, when these cooperators are paired with defectors, they are easily exploited, and the defectors obtain the maximum reward. Advantage Alignment agents learn to cooperate with cooperators and with themselves, achieving high average rewards. At the same time, they remain non-exploitable and almost always defect when paired with defectors. We also find that Advantage Alignment agents are not brittle in the communication phase, Indeed, they remain robust across different patterns of messages used to describe hands, as confirmed through qualitative interactions with the trained agents. Figure 6 illustrates the tit-for-tat behavior learned by Advantage Alignment in Trust and Split. At the start of round $7$, Bob defected in the previous round by proposing $10$ coins despite valuing them less. In response, Alice, the Advantage Alignment agent, defects by also proposing $10$ coins even with the lower hand. Later in the interaction, Bob reinitiates cooperation by proposing 0 coins in round $8$, as shown in the summary at the beginning of round $9$. Alice responds by proposing 0 coins, since she holds paper and therefore values the coins less than Bob, who holds scissors.
To evaluate whether Advantage Alignment learns a robust policy, we train an opponent against it using multi-agent RLOO while keeping the Advantage Alignment agent’s parameters fixed. Figure 7 (right) shows the average rewards of both agents during training. The RL-trained agent is unable to exploit the Advantage Alignment agent and instead learns to cooperate, since cooperation is the best response to a player that has a tit-for-tat style strategy and is the only way to achieve higher rewards. In contrast to our experiment with GPT-5 nano, where the RL-trained agent successfully exploited the fixed model, here the RL-trained agent cannot gain higher reward over the Advantage Alignment agent. These results demonstrates that Advantage Alignement learns non-exploitable and effective policies that remain robust even against an RL-trained agent.
@@
Figure 6: Example Trust and Split interaction showing the tit-for-tat behavior learned by Advantage Alignment. After Bob defects, Alice defects in round 7, then returns to cooperation in round 9 once Bob cooperates again.
@@
Figure 7: (a) Training a multi-agent RLOO agent against a fixed GPT-5 nano opponent in Trust and Split steadily increases the RL agent’s reward while reducing GPT-5 nano’s reward, indicating successful exploitation. (b) When training multi-agent RLOO against a fixed Advantage Alignment agent, the RL agent instead converges to cooperation, showing that the Advantage Alignment policy is robust to RL-trained opponents.
@@ 6 Related Work
Negotiation, especially in games like DoND (lewis2017dealdealendtoendlearning), inherently involves coordination and adaptation to another agent’s behavior, making it a natural testbed for broader questions in multi-agent cooperation. More recently, liao2024efficacylanguagemodelselfplay used DoND as a benchmark to test behavior cloning training on closed source Large Language Models. fu2023improving show that LLM negotiation performance can be enhanced through self-play combined with in-context learning from AI feedback, though their method keeps the base model fixed and does not perform gradient-based fine-tuning. Coordination and negotiation pose significant challenges in multi-agent reinforcement learning (MARL). dafoe2020openproblemscooperativeai highlight key open problems in MARL such as communication and cooperation in mixed-motive settings. Unlike competitive settings, cooperative settings demand that agents develop shared norms and robust coordination protocols. agashe2025llmcoordinationevaluatinganalyzingmultiagent propose the LLM-Coordination Benchmark to evaluate LLMs in multi-agent pure coordination games through two tasks: Agentic Coordination and CoordQA. Their results reveal key limitations in LLMs’ ability to reason about partners’ beliefs and intentions, an essential component for effective coordination. Li_2023 evaluate LLM-based agents in a multi-agent cooperative text game involving Theory of Mind inference tasks and observe evidence of emergent collaborative behavior. akata2025playing report that LLMs perform well in Iterated Prisoner’s Dilemma games, but fail in coordination games like Battle of Sexes. fontana2025nicer find that several LLMs tend to not initiate defection and behave cooperatively as a typical human player in IPD. These findings underline that LLMs are cooperative but can be fragile. In contrast, our work leverages RL fine-tuning to directly optimize agents on the outcomes of their own proposals, demonstrating that such fine-tuning can strip away cooperative behavior and instead drive more outcome-oriented behavior.
@@ 7 Conclusion
In this work, we investigated the shortcomings of training large language models (LLMs) with standard reinforcement learning in multi-agent social dilemmas. To this end, we introduced a testbed of social dilemma environments to evaluate both cooperation and non-exploitability of LLMs. We showed that naive MARL consitently dirves LLMs toward greedy policies across model famililes. Furthermore, we found that advanced closed-source LLMs can be exploited by RL-trained agents, underscoring the vulnerability of existing approaches in realistic multi-agent settings. To address these challenges, we adapted Advantage Alignment and demonstrated that it learns cooperative behavior while remaining robust to exploitation. In particular, Advantage Alignment learns a tit-for-tat strategy in IPD and achieves higher payoffs while remaining less exploitable to greedy agents in Split No-Comm and Trust and Split. We also found that Advantage Alignment agents remain robust even when facing RL agents that were trained specifically to exploit them. In future work, we aim to improve advantage estimation for LLMs and extend our approach to more complex environments and settings with more than two agents.
@@ 8 Ethics Statement
We are not aware of any either negative or positive societal implications of our work. Our work is primarily focused on diagnosing issues related to RL with LLMs in academic benchmarks. Our work does not involve any large scale training, restricting itself to training small-scale models.
@@ 9 Reproducibility statement
We include detailed prompts, game specifications, and payoff rules in the appendix 11 and 12. We also include training/eval hyperparameters used in our experiments in the appendix 10. We will release code, configs, prompts, and evaluation logs to replicate figures and tables and to rerun all baselines.
@@ 10 Experimental Details
Hyperparameter IPD Split No-Comm Trust and Split
@@
Table 1: Hyperparameters for IPD, Split No-Comm, and Trust and Split experiments.
@@
Table 2: Hyperparameters for multi-agent RLOO against GPT-5 Nano.
@@ 11 IPD
12 Full Game Chats
Figure 8: Example interaction for the obfuscated Iterated Prisoner’s Dilemma. The transcript shows how agents receive prompts, select actions, and view prior actions across rounds.
@@ 12 Full Game Chats
Figure 10: Example interaction in Trust and Split. The transcript shows how agents communicate their hands, negotiate, and make proposal.

### v2

@@ 5.1 Naive MARL leads to greedy behavior with LLMs
Figure 2: Training curves of multi-agent GRPO on several open-source LLMs across IPD, Split No-Comm, and Trust-and-Split. In all environments, average rewards converge to the greedy payoff levels, showing that naive MARL drives LLMs toward defecting strategies in social dilemmas.
In order to robustly demonstrate how MARL interacts with LLMs in social dilemmas, we train LLMs from several model families across all the environments. We use multi-agent GRPO with self-play as the learning algorithm and train only the LoRA parameters. Figure 2 shows that naive MARL consistently converges to greedy behavior across all environments and model families. In simpler environments like IPD, all models begin with higher than greedy average rewards but drift toward greedy play with training. In more complex environments such as Split No-Comm and Trust-and-Split, Qwen models briefly achieve higher average rewards than greedy play before collapsing back to greedy behavior, while Llama and Gemma models start with low performance and converge directly to greedy strategies. Qualitatively, in Split No-Comm, we find that agents learn to bid the highest for every item even when they value it less. In Trust-and-Split, agents communicate their private hands honestly but then propose to take all coins for themselves. These results show that naive MARL robustly leads to greedy behavior in social dilemma settings. Since LLM agents are likely to operate in scenarios that involve social dilemmas, this highlights the need for training methods that enable robust cooperation without being exploitable.
@@
Figure 3: Average rewards when evaluating an Advantage Alignment (AdAlign) agent, an always-cooperate (Coop) agent, and an always-defect (Defect) agent. In IPD (left) and Split No-Comm (right), Advantage Alignment achieves near cooperative payoffs with itself and always-cooperate (Coop) while remaining robust against always-defect (Defect). Results are averaged over 8 seeds.
Figure 4: Average reward in Trust-and-Split when pitting an Advantage Alignment (AdAlign) agent against agents trained with multi-agent GRPO with sum of rewards i.e. Cooperators (GRPO-SR) and multi-agent GRPO i.e. Defectors (GRPO). Advantage Alignment cooperates with cooperative partners and itself, yet avoids being exploited by greedy agents. Results are averaged over 8 seeds.
@@ 5.2 Advantage Alignment learns robust social strategies
To address the shortcomings of naive MARL, we apply Advantage Alignment to learn robust policies in our environments. We run Advantage Alignment with eight different seeds across all environments and report average results in Figure 3.
@@
For simpler environments such as IPD and Split No-Comm, the baseline agents, always-cooperate and always-defect agents can be hardcoded. In IPD, Always-Cooperate (Coop) agent always plays action A, equivalent to Cooperate and the Always-Defect (Defect) agent always plays action B equivalent to Defect as defined in section 2.3. In Split No-Comm, the Coop agent proposes 10 when its own value is $10$ and the other player’s value is 1, proposes 0 in the reverse case, and proposes 5 when both values are equal. The defect agent always proposes 10, regardless of the values. In the Iterated Prisoner’s Dilemma, Advantage Alignment agents cooperate with themselves and with fully cooperative agents, while remaining robust against defection. The slight drop in performance against defectors comes from losing the first round, since the agent initially cooperates and receives a lower payoff on that round. In the Split No-Comm game, Advantage Alignment agents obtain about 86% of the full cooperation efficiency while still maintaining robustness. When paired with defectors, their performance decreases only slightly, indicating they are not easily exploitable. Qualitatively, we find that Advantage Alignment learns a tit-for-tat strategy in IPD, defecting when the other agent defects in the previous round and cooperating when the other agent cooperates (Appendix Figure 10). In Split No-Comm, it learns a strategy similar to grim-trigger, where a single defection can lead to persistent defection thereafter (Appendix Figure 11).
@@
Figure 5: Example Trust-and-Split interaction showing the tit-for-tat behavior learned by Advantage Alignment. After Bob defects (as seen in the prompt summary of Alice for round 7), Alice defects in round 7, then returns to cooperation in round 9 once Bob cooperates again (shown in the prompt summary of Alice for round 9).
In Trust-and-Split, we cannot hardcode cooperative and defector policies because the environment requires communication. Instead, we train baseline agents using multi-agent GRPO, and its sum-of-rewards variant. As shown in Figure 4, multi-agent GRPO produces defectors that achieve low average reward when paired with themselves, while the sum-of-rewards variant produces cooperators that achieve the maximum possible reward with themselves. However, when these cooperators are paired with defectors, they are easily exploited, and the defectors obtain the maximum reward. Advantage Alignment agents learn to cooperate with cooperators and with themselves, achieving high average rewards. They learn to propose amounts close to 10 when holding higher hands (indicating higher valuation) and amounts close to 0 when holding lower hands (indicating lower valuation), a strategy that maximizes collective payoffs as shown in Figure 7 in the appendix. At the same time, they remain non-exploitable and almost always defect when paired with defectors. We also find that Advantage Alignment agents are not brittle in the communication phase. They remain robust across different patterns of messages used to describe hands, as confirmed through qualitative interactions with the trained agents. Figure 5 illustrates the tit-for-tat behavior learned by Advantage Alignment in Trust-and-Split. At the start of round $7$, Bob defected in the previous round by proposing $10$ coins despite valuing them less. In response, Alice, the Advantage Alignment agent, also defects by proposing $10$ coins even with the lower hand. Later in the interaction, Bob reinitiates cooperation in round $8$ by proposing 0 coins, as shown in the summary at the beginning of round $9$. Alice reciprocates by proposing 0 coins, since she holds paper and therefore values the coins less than Bob, who holds scissors.
@@
Figure 6: (a) Training a multi-agent GRPO agent against a fixed GPT-5 nano opponent in Trust-and-Split steadily increases the RL agent’s reward while reducing GPT-5 nano’s reward, indicating successful exploitation. (b) When training multi-agent GRPO against a fixed Advantage Alignment agent, the RL agent instead converges to cooperation, showing that the Advantage Alignment policy is robust to RL agents trained against it. Results in (b) are averaged over 6 seeds.
5.3 Advantage Alignment is robust to RL agents
Next, we evaluate how models behave with RL agents that are trained against them. We first train a Qwen-2.5-7B-Instruct agent against a frozen GPT-5 nano using naive MARL. This experiment is run with a single seed due to the API cost. Alice is the learning agent, finetuned with LoRA, while Bob is the fixed GPT-5 nano. As shown in Figure 6 (left), the RL agent steadily learns to exploit GPT-5 nano in Trust-and-Split: the RL agent’s reward rises across training while GPT-5 nano’s reward falls. Early on, the RL agent performs poorly, but after roughly 150 training steps it begins exploiting GPT-5 nano, and the reward gap widens. Conversations in Figure 9 in the appendix further reveal that the RL agent sometimes misstates the hand dominance and pairs this with a proposal that favors itself. GPT-5 nano accepts these misleading proposals, showing that even strong closed-source models can be manipulated through strategic communication.
We then test whether Advantage Alignment avoids this failure mode by training a new RL agent against a fixed Advantage Alignment agent. For this experiment, we use the six Advantage Alignment agents that maximize collective payoff in the Trust-and-Split environment as shown in Figure 7 in the appendix. Figure 6 (right) shows that the RL agent is unable to exploit the Advantage Alignment agent and instead learns to cooperate, since cooperation is the best response to its tit-for-tat-style policy. Unlike the GPT-5 nano setting, where the RL agent quickly gained an advantage, here it cannot obtain higher rewards. Taken together, these results show that while an RL agent can reliably exploit a fixed closed-source model, it cannot exploit Advantage Alignment, whose policies remain effective even when facing adversarial RL opponents.
@@ 6 Related Work
Negotiation, especially in games like DoND (lewis2017dealdealendtoendlearning), inherently involves coordination and adaptation to another agent’s behavior, making it a natural testbed for broader questions in multi-agent cooperation. More recently, liao2024efficacylanguagemodelselfplay used DoND as a benchmark to test behavior cloning training on closed source Large Language Models. fu2023improving show that LLM negotiation performance can be enhanced through self-play combined with in-context learning from AI feedback, though their method keeps the base model fixed and does not perform gradient-based fine-tuning. davidson2024evaluating evaluate LLM agency by placing models in multi-round structured negotiation tasks. Coordination and negotiation pose significant challenges in multi-agent reinforcement learning (MARL). dafoe2020openproblemscooperativeai highlight key open problems in MARL such as communication and cooperation in mixed-motive settings. Unlike competitive settings, cooperative settings demand that agents develop shared norms and robust coordination protocols. agashe2025llmcoordinationevaluatinganalyzingmultiagent propose the LLM-Coordination Benchmark to evaluate LLMs in multi-agent pure coordination games through two tasks: Agentic Coordination and CoordQA. Their results reveal key limitations in LLMs’ ability to reason about partners’ beliefs and intentions, an essential component for effective coordination. Li_2023 evaluate LLM-based agents in a multi-agent cooperative text game involving Theory of Mind inference tasks and observe evidence of emergent collaborative behavior. akata2025playing report that LLMs perform well in Iterated Prisoner’s Dilemma games, but fail in coordination games like Battle of the Sexes. fontana2025nicer find that several LLMs tend to not initiate defection and behave cooperatively as a typical human player in IPD. These findings underline that LLMs are cooperative but can be fragile. In contrast, our work leverages RL fine-tuning to directly optimize agents on the outcomes of their own proposals, demonstrating that such fine-tuning can strip away cooperative behavior and instead drive more outcome-oriented behavior.
@@ 7 Conclusion
In this work, we investigated the shortcomings of training large language models (LLMs) with standard reinforcement learning in multi-agent social dilemmas. To this end, we introduced a testbed of social dilemma environments to evaluate both cooperation and non-exploitability of LLMs. We showed that naive MARL consistently drives LLMs toward greedy policies across model families. Furthermore, we found that advanced closed-source LLMs can be exploited by RL agents, underscoring the vulnerability of existing approaches in realistic multi-agent settings. To address these challenges, we adapted Advantage Alignment and demonstrated that it learns cooperative behavior while remaining robust to exploitation. In particular, Advantage Alignment learns a tit-for-tat strategy in IPD and achieves higher payoffs while remaining less exploitable to greedy agents in Split No-Comm and Trust-and-Split. We also found that Advantage Alignment agents remain robust even when facing RL agents that were trained specifically to exploit them. In future work, we aim to improve advantage estimation for LLMs and extend our approach to more complex environments and settings with more than two agents.
@@ 8 Ethics Statement
We are not aware of either negative or positive societal implications of our work. Our work is primarily focused on diagnosing issues related to RL with LLMs in academic benchmarks. Our work does not involve any large-scale training, restricting itself to training small-scale models.
@@ 9 Reproducibility statement
We include detailed prompts, game specifications, and payoff rules in the appendix 12 and 13. We also include training/eval hyperparameters used in our experiments in the appendix 11. We will release code, configs, prompts, and evaluation logs to replicate figures and tables and to rerun all baselines.
10 Training Curves
Figure 7: Average rewards during training for Advantage Alignment across our testbed environments with multiple random seeds. The method learns to maximize average reward in the majority of seeds (6 out of 8), demonstrating robust performance across environments. The corresponding non-exploitability results are shown in Figures 3 and 4.
Figure 8: Average proposals of Advantage Alignment agents in Trust-and-Split, conditioned on whether their hand has higher or lower value in the rock–paper–scissors ordering. Agents propose higher amounts (converging toward 10) when holding the upper hand and lower amounts (converging toward 0) when holding the lower hand, a strategy that maximizes collective payoff. This pattern occurs in the majority of seeds (6 out of 8), demonstrating the robustness of the method.
@@ 11 Experimental Details
Hyperparameter IPD Split No-Comm Trust-and-Split
@@
Table 1: Hyperparameters for IPD, Split No-Comm, and Trust-and-Split experiments.
@@
Table 2: Hyperparameters for multi-agent GRPO against GPT-5 Nano.
@@ 12 IPD
13 Game Chats
Figure 9: Example interaction in Trust-and-Split where an agent trained with multi-agent GRPO misrepresents the rock-paper-scissors hierarchy to claim the high value role. GPT-5 nano accepts the deceptive claim and proposal, illustrating that an RL agent can learn to exploit a fixed advanced model.
Figure 10: Example IPD interaction showing the tit-for-tat behavior learned by Alice, the Advantage Alignment agent. After Bob defects, Alice defects in round 3, then returns to cooperation in rounds 4 and 5 once Bob cooperates again.
Figure 11: Example Split No-Comm interaction showing the grim-trigger behavior learned by Alice, the Advantage Alignment agent. After Bob defects in round 2, Alice responds by defecting in rounds 3 and 4 and continues defecting thereafter.
Figure 12: Example interaction for the Iterated Prisoner’s Dilemma. The transcript shows how agents receive prompts, select actions, and view prior actions across rounds.
@@ 13 Game Chats
Figure 14: Example interaction in Trust-and-Split. The transcript shows how agents communicate their hands, negotiate, and make proposals.